# TG-NAS: Leveraging Zero-Cost Proxies with Transformer and Graph Convolution Networks for Efficient Neural Architecture Search

### v1

@@ 3.3 GCN Proxy Trainer
After completing the universal architecture encoding and operator feature embedding, we employ a two-layer graph convolution network (GCN) as our prediction model. With normalization trick, GCN can be defined as
$H=X*_{G}G_{\Theta}=f(\bar{A}X\Theta),$ (2)
@@
where $\bar{A}=\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}$, $\tilde{A}=A+I_{n}$, and $\tilde{D_{ii}}=\sum_{j}\tilde{A}_{ij}$. To prevent overfitting to a particular training space, we incorporate graph normalization [5] and weight decay techniques. Our overarching objective is to deliver a universally applicable pretrained predictor model that requires no tuning for new search spaces. Consequently, the GCN predictor model is subject to heavy regularization. An additional crucial factor influencing our choice of GCN over other prediction models is its capability to handle vast differences in architectures. Given the varying dimensions of the adjacency matrix (from unseen search spaces) and operator matrix (from unseen operators), GCN emerges as a suitable choice, demonstrating flexibility in accommodating diverse architectural structures.
Figure 2: GCN Predictor Functional Validation on NAS-Bench-101
To assess the predictor’s applicability, we partition the NAS-Bench-101 space into train/validation splits ranging from 90% to 1%. The results, depicted in Fig. 2, demonstrate the predictor’s effectiveness, even when trained on only 1% of the architecture data. This is evident in the strong performance indicated by Kendall’s $\tau$ and Spearman’s $\rho$ correlation coefficients. Exploring the use of different graph neural networks (GNNs) such as graph attention network (GAT) [43] may yield different results and behaviors. We consider this as an open area for further investigation.
3.4 Pruning-Evolutionary Hybrid Searching
The search strategy generally falls into two types. The first type involves applying black-box optimization techniques (reinforcement learning and evolution search etc.) in conjunction with a sampling-based method to explore the predefined search space. These methods sample and evaluate architectures to optimize the searching agent’s reward. The other type is close to one-shot techniques, where a hypernetwork will be introduced with architecture representation relaxation to enable searching using gradient descent. It can be applied to all DAG-based search spaces [28] due to its continuous hyperparameter for the architecture. However, when considering the cell-based search space, the computational burden of these sampling-based one-shot methods remains dependent on the scale of the search space and hence it has limited effectiveness. Consider a network comprising interconnected cells in a directed acyclic graph structure. Each cell contains $E$ edges, with each edge selecting a single operator from a set of $\lvert O\rvert$ potential candidates. Consequently, there are $|O|^{E}$ distinct cells, and during the sampling-based search process, $\alpha|O|^{E}$ networks must be sampled. The parameter $\alpha$ represents the sampling process efficiency, where a smaller value can help find superior architectures faster.
Meanwhile, evolutionary and genetic algorithms have been commonly used to optimize the NAS[46]. To enhance the efficiency of the search, we propose a pruning-evolutionary hybrid searching method. The pruning method, inspired by [19], reduces the search cost from $\alpha|O|^{E}$ to $|O|\cdot E$. Algorithm 1 shows our proposed algorithm.
@@ 3.4 Pruning-Evolutionary Hybrid Searching
For the pruning search part, we start with a supernetwork $N_{init}$ encompassing all possible operators for each edge. Then we iteratively prune one operator on each edge until the current supernet $N_{t}$ transforms into a single-path network, representing the final searched architecture. For the evolutionary search part, we first initialize the entire population of architecture candidates with continuous parameters, where we map all operators evenly into the value between 0 and 1. Then mutation operation and crossover operation are performed to produce a new child. After generating all the offspring, the selection process will kick in using our proposed proxy to select the elite candidates for the next generation. After finishing both parts of searching, we select the final architecture from the best of the two search results. This hybrid searching method is valid and efficient due to the effectiveness of the proposed zero-cost proxy.
Data: $M_{0}$ supernet, $E$ edges in each cell, every edge have $O$ operators
while $M_{t}$ is not a single path network do
       for $Cell\_type$, $ct$ in $M_{t}$ do
             for edge $E_{i}$, operator $O_{j}$ in $ct$, $E_{i}$ do
                   for operator $O_{j}$ in $E_{i}$ do
                         $P_{t,e_{i},o_{j}}\leftarrow GCN\_Inference(M_{t,e_{i},o_{j}})$;
                         $\Delta P_{t,e_{i},o_{j}}\leftarrow P_{t,e_{i},o_{j}}-P_{t,e_{i}\backslash o_{j}}$;
                   end for
             end for
            Rank $\leftarrow$ index of $o_{j}$ in descending of $\Delta P_{t}$;
       end for
      for $e_{i}$ in $E$ do
             $o_{p}\leftarrow argmin\{S(o_{p})\>o_{p}\in e_{i}\}$;
             $M_{t+1}\leftarrow M_{t}-o_{p}$;
       end for
      $t\leftarrow t+1$;
end while
Data: $NP$ population size
$g\leftarrow 0$;
while $|pop|<NP$ do
       $pop_{i}\leftarrow random\_configuration()$;
       ${pop^{\prime}}_{i}\leftarrow discretized\_architecture(pop_{i})$;
       $fitness_{i}\leftarrow GCN\_Inference({pop^{\prime}}_{i})$;
end while
while $g<g_{max}$ do
       $V_{g}\leftarrow mutate(pop_{g})$;
       $U_{g}\leftarrow crossover(V_{g},pop_{g})$;
       ${U^{\prime}}_{g}\leftarrow discretized\_population(U_{g})$;
       $fitness_{g}\leftarrow GCN\_Inference({U^{\prime}}_{g})$;
       $pop_{g+1},fitness_{g+1}\leftarrow select(pop_{g},U_{g})$;
end while
return $Max(M_{t},POP_{g})$;
Algorithm 1 TG-NAS Pruning-Evolutionary Hybrid Search Algorithm
@@ 4 Experiment and Results Analysis
4.1 Independent Analysis and Combination with other proxies
@@
Figure 3: Spearman’s $\rho$ Correlation for All Pairs of ZC Proxies on NAS-Bench-201
@@
Multiple proxies could be combined to enhance the performance prediction accuracy. However, certain proxies could be highly correlated and do not provide extra information. Therefore, not all combinations provide good performance. We conducted assessments of zero-cost proxies on NAS-Bench-201 search spaces using the aforementioned three datasets. Additionally, by evaluating the same search space across different tasks, we aimed to determine if data- and task-independent zero-cost proxies offer universally applicable rankings. Our evaluation encompassed 14 zero-cost proxies, including our TG-NAS proxy result. We randomly sampled 1000 architectures from the search space and evaluated each zero-cost proxy metric for each architecture. Subsequently, we computed the $Spearman\textquoteright s$ rank correlation between each proxy.
@@
Our analysis revealed a trend where the rank correlations of some zero-cost proxies were highly correlated, such as $flops$ (number of floating-point operations) and $params$ (model parameter count), which is expected due to the relationship between parameters and computation in deep learning. This prompted us to compute the full correlations between all 14 pairs of zero-cost proxies and analyze their behaviors.
@@
Fig. 10 shows the heatmap of correlations between pairs of popular proxies (calculated with CIFAR-10 results; CIFAR-100 and ImageNet16-120 results are available in the supplementary materials). We observed a consistent trend where $flops$ and $params$ exhibited high correlations with each other. Additionally, $synflow$ [1] often showed high correlations with $flops$ and $params$, consistent with recent work by Ning et al. [32], which demonstrated that $synflow^{\prime}s$ value increases with the number of parameters in the architecture. We also noted high correlations between $grad\_norm$[1], $snip$[18], $grasp$[44], $nwot$[31], and $fisher$[42], occasionally with $l2\_norm$[1] as well, as they all leverage gradient saliency matrix information. Moreover, $epe\_nas$[29] and jacobian covariance ($jacov$)[31] were highly correlated with each other.
@@
Table 3: Kendall’s $\tau$ and Spearman’s $\rho$ correlation between various zero-cost proxies
Proxy Name Params FLOPs SNIP Fisher Synflow Zen-score grad-norm TG-NAS (ours) Kendall’s $\tau$ 0.55 0.54 0.41 0.22 0.54 0.29 0.37 0.57 Spearman’s $\rho$ 0.74 0.73 0.58 0.36 0.73 0.38 0.54 0.77
4.2 Result Analysis of CIFAR-10 on NAS-Bench-201 Space
(a) Ground Truth vs. TG-NAS Pridiction
(b) FLOPs vs. CIFAR-10 Accuracy
Figure 4: TG-NAS Correlation Evaluation with Ground Truth on NAS-Bench-201 Space
For testing on NAS-Bench-201, our predictor model was trained using the entire NAS-Bench-101 architecture set, and the obtained predictions were employed to guide the search on the new space, NAS-Bench-201, using the proposed method. The Kendall’s $\tau$ and Spearman’s $\rho$ correlation scores displayed in Table 3 indicate that our TG proxy outperforms other zero-cost proxies when evaluated on NAS-Bench-201 spaces. Fig. 4 demonstrates that our TG proxy is highly positively correlated with the architecture’s test accuracy. Additionally, as illustrated in Table 4, our TG-NAS outperforms the majority of zero-shot NAS approaches and is comparable to the state-of-the-art TE-NAS results, achieving 93.75% top-1 accuracy on CIFAR-10, while consuming only a fraction of the searching time of prior works.
Notably, despite the efficiency claims of zero-cost NAS methods over conventional NAS due to avoiding the training of sampled architectures, variations in computational cost and search time persist among them. For example, TE-NAS [4] required over 4 GPU hours, while ZiCo [21] demanded over 10 GPU hours for the search on ImageNet on the DARTS space. This discrepancy arises from the fact that several proxy computations necessitate at least one forward pass or a forward-backward pass, requiring multiple runs for result stabilization. Additionally, many are data-dependent, limiting their general applicability and extending search times.
In contrast, TG-NAS completed the search in about 40 seconds, achieving up to a 300×300$\times$300 × speedup compared to other zero-cost methods due to the lightweight nature of the proposed predictor model. Notably, the comparison of search times does not factor in the predictor model training time (the model training required 1.5 GPU hours), as it is a one-time training effort, and the pretrained model can be distributed for uses in various scenarios and search spaces.
@@ 4.2 Result Analysis of CIFAR-10 on NAS-Bench-201 Space
4.3 Result Analysis of ImageNet on DARTS Space
For the DARTS search space, our GCN prediction model undergoes a slightly different training process, incorporating additional data from NAS-Bench-201 into the training set. Consequently, our training dataset comprises a total of 423,000 NAS-Bench-101 architectures and an additional 15,625 architectures from NAS-Bench-201. The labels for the NAS-Bench-101 portion are derived from the CIFAR-10 evaluation accuracy. However, the NAS-Bench-201 labels differ slightly due to additional CIFAR-100 and ImageNet-16-120 results. To mitigate bias arising from the monolithic dataset, we amalgamate the normalized accuracies of CIFAR-10, CIFAR-100, and ImageNet-16-120 to construct a unified ground truth label for NAS-Bench-201 architectures. Indeed, while introducing an additional proxy is necessary due to the absence of actual results on the ImageNet dataset in those NAS benchmarks, it’s worth noting that CIFAR-10/100 performance often serves as a reliable indicator for ImageNet performance. The final discovered cell architecture can be seen in Fig. 5. After the target cell was determined, the final network was constructing by stacking 14 cells, with the initial channel number set to 48. Performance results are shown in Table 5. Our approach achieves a top-1/5 accuracy of 74.5% and 91.9% respectively, comparable to other works on ImageNet. Notably, TG-NAS only took less than 2 minutes of search time on one NVIDIA RTX 4090 GPU. We expect that performance can be further enhanced if incorporating additional training samples from supplementary NAS benchmarks.
Table 5: Comparison with Recent NAS Works on ImageNet with the mobile setting.
Name of Works Test Accuracy (%) Search Cost (GPU Days) Params (M) Search Method Top-1 Top-5 PNAS [26] 74.2 91.9 225 5.1 Bayesian Optimization AmoebaNet-C [36] 75.7 92.4 3150 6.4 Evolution NASNet-A [55] 74.0 91.6 2000+ 5.3 Reinforcement Learning DARTS [28] 73.3 91.3 4.0 4.7 Gradient-based SNAS [49] 72.7 91.8 1.5 4.3 Gradient-based BayesNAS [54] 73.5 91.1 0.2 3.9 Gradient-based ProxylessNAS [3] 75.1 92.5 8.3 7.1 Gradient-based TE-NAS [7] 74.2 92.2 0.17 5.4 Theoretical analysis BONAS-B [40] 75.2 92.7 5.0 4.5 Bayesian Optimization PC-DARTS [51] 74.9 92.2 0.1 5.3 Gradient-based TG-NAS (ours) 74.5 91.9 0.0014 5.6 Model-based Predictor
@@ 5 Conclusion
In this work, We present TG-NAS, a model-based predictor framework that is generally applicable for new search spaces with unseen new operators. It integrates a transformer-based operator embedding generator with a graph convolution network and functions as a ZC proxy. TG-NAS offers advantages in terms of data independence, cost-effectiveness, and generality. Our experiments demonstrate its superiority over existing proxies across various NAS benchmarks, positioning it as a foundational element for efficient architecture search. TG-NAS achieves up to a 300$\times$ improvement in search efficiency compared to previous state-of-the-art zero-cost proxy methods. Notably, it discovers competitive models with 93.75% CIFAR-10 accuracy on the NAS-Bench-201 space and 74.5% ImageNet top-1 accuracy on the DARTS space.
(a) Normal Cell
(b) Reduction Cell
Figure 5: Discovered DNN Cell Architecture on the DARTS Space

### v2

@@ 3.3 GCN Proxy Predictor
After completing the universal architecture encoding and operator feature embedding, we employ a three-layer graph convolution network (GCN) [11] as our prediction model. With the normalization trick, GCN can be defined as
$H=X*_{G}G_{\Theta}=f(\bar{A}X\Theta)$ (2)
@@
where $\bar{A}=\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}$, $\tilde{A}=A+I_{n}$, and $\tilde{D_{ii}}=\sum_{j}\tilde{A}_{ij}$. To prevent overfitting to a particular training search space, we incorporate graph normalization and weight decay techniques. Our overarching objective is to deliver a universally applicable pre-trained predictor model that requires no tuning for new search spaces. Consequently, the GCN predictor model is subject to heavy regularization. An additional crucial factor influencing our choice of GCN over other prediction models is its capability to handle vast differences in architectures. Given the varying dimensions of the adjacency matrix (from unseen search spaces) and operator matrix (from unseen operators), GCN emerges as a suitable choice, demonstrating flexibility in accommodating diverse architectural structures.
3.4 Evolutionary Searching
@@
Evolutionary and genetic algorithms have been commonly used to optimize the NAS [44]. To enhance the efficiency of the search, we adopt similar approach. We first initialize the entire population of architecture candidates with continuous parameters, where we map all operators evenly into the value between 0 and 1. Then mutation operation and crossover operation are performed to produce a new child. After generating all the offspring, the selection process will kick in using our proposed proxy to select the elite candidates for the next generation. In the end, we select the final returned architecture as the search result.
Figure 3: Spearman’s $\rho$ rank between ZC proxy values and ground truth accuracies, for 14 ZC proxies and across 12 NAS benchmarks (the higher the better).
@@ 4 Experiment and Results Analysis
In our evaluation, we tested 12 NAS benchmarks similar to NAS-Bench-Suite-Zero [12], including NAS-Bench-201 (CIFAR-10, CIFAR-100, and ImageNet16-120) [5], NAS-Bench-301 (CIFAR-10) [52], and TransNAS-Bench-101 Micro (Jigsaw, Object Classification, Scene Classification, Autoencoder, Room Layout, Surface Normal, and Semantic Segmentation) [6]. We excluded NAS-Bench-101 [51] from the comparison because it was used as training data for our GCN predictor. To prevent information leakage and ensure a fair comparison, we removed it from subsequent evaluations.
For the GCN predictor, we employed a three-layer GCN with a weight decay of $3\times 10^{-6}$ and an initial learning rate of 0.01, using the ReduceLROnPlateau scheduler and the AdamW optimizer. The model was trained for 150 epochs to ensure convergence. To minimize training cost and highlight the effectiveness of our novel architecture and operator encoding method in generalizing to unseen operators, the GCN was trained solely on NAS-Bench-101 (CIFAR-10) [51].
4.1 Generalizability Analysis of ZC Proxies Across 12 Benchmarks
In Figure 3, we report the Spearman’s $\rho$ correlation between each zero-cost (ZC) proxy and the ground-truth performance for each benchmark. Among existing proxies, Jacov[28] and Zen Score[21] exhibit the highest rank correlations across all benchmarks. However, our proposed TG-proxy consistently outperforms them, achieving state of the art result.
Notably, in challenging tasks such as the TransNAS-Bench-101-Micro [6] Autoencoder and Room Layout, where most ZC proxies perform poorly, TG-proxy maintains strong predictive power. On the widely used NAS-Bench-201[5] benchmarks, although several proxies demonstrate good performance, TG-proxy still ranks the highest. While methods such as Snip[14] and Grasp[42] perform competitively on NAS-Bench-201, they generalize poorly to other benchmarks and are even outperformed by simple heuristics like Params and FLOPs.
Overall, these results highlight that TG-proxy not only achieves state-of-the-art performance on average accrossing search spaces but also with the lowest standard divination across diverse NAS benchmarks that demonstrates superior generalizability.
4.2 Independent Analysis of proxies
@@
Figure 4: Spearman’s $\rho$ Correlation between ZC proxy scores of CIFAR-10 on NAS-Bench-201.
@@
Multiple proxies could be combined to enhance the performance prediction accuracy. However, certain proxies could be highly correlated and do not provide extra information. Therefore, not all combinations provide good performance. We conducted assessments of zero-cost proxies on NAS-Bench-201 search spaces using the aforementioned three datasets. Additionally, by evaluating the same search space across different tasks, we aimed to determine if data- and task-independent zero-cost proxies offer universally applicable rankings. Our evaluation encompassed 14 zero-cost proxies, including our TG-NAS proxy result. We randomly sampled 1000 architectures from the search space and evaluated each zero-cost proxy metric for each architecture. Subsequently, we computed the Spearman’s rank correlation between each proxy.
@@
Our analysis revealed a trend where the rank correlations of some zero-cost proxies were highly correlated, such as $FLOPs$ and $params$ , which is expected due to the relationship between parameters and computation in deep learning. This prompted us to compute the full correlations between all 14 pairs of zero-cost proxies and analyze their behaviors.
@@
Figure 8 shows the heatmap of correlations between pairs of popular proxies (calculated with CIFAR-10 results; CIFAR-100 and ImageNet16-120 results are available in the supplementary materials). We observed a consistent trend where $FLOPs$ and $params$ exhibited high correlations with each other. Additionally, $synflow$ [1] often showed high correlations with $FLOPs$ and $params$, consistent with recent work by Ning et al. [29], which demonstrated that $synflow^{\prime}s$ value increases with the number of parameters in the architecture. We also noted high correlations between $grad\_norm$ [1], $snip$ [14], $grasp$ [42], $nwot$ [28], and $fisher$ [41], occasionally with $l2\_norm$ [1] as well, as they all leverage gradient saliency matrix information. Moreover, $epe\_nas$ [26] and jacobian covariance ($jacov$) [28] were highly correlated with each other.
@@
4.3 NAS Result on NAS-Bench-201
(a) CIFAR-10
(b) CIFAR-100
(c) ImageNet16-120
Figure 5: TG-NAS vs. Ground Truth ranking on NAS-Bench-201 Space
Figure 7 demonstrates that our TG proxy is highly positively correlated with the architecture’s accuracy ranking. Additionally, as illustrated in Table 3, our TG-NAS outperforms the majority of zero-shot NAS approaches and is comparable to the state-of-the-art TE-NAS results, achieving 93.75% top-1 accuracy on CIFAR-10, while consuming only a fraction of the searching time of prior works.
@@ 4.3 NAS Result on NAS-Bench-201
Table 4: Comparison with Recent NAS Works on ImageNet with the Mobile Setting
Name of Works Test Accuracy (%) Search Cost (GPU Days) Params (M) Search Method Top-1 Top-5 PNAS [23] 74.2 91.9 225 5.1 Bayesian Optimization AmoebaNet-C [35] 75.7 92.4 3150 6.4 Evolution NASNet-A [56] 74.0 91.6 2000+ 5.3 Reinforcement Learning DARTS [25] 73.3 91.3 4.0 4.7 Gradient-based SNAS [47] 72.7 91.8 1.5 4.3 Gradient-based BayesNAS [54] 73.5 91.1 0.2 3.9 Gradient-based ProxylessNAS [2] 75.1 92.5 8.3 7.1 Gradient-based TE-NAS [4] 73.8 91.7 0.05 6.3 Theoretical analysis BONAS-B [39] 75.2 92.7 5.0 4.5 Bayesian Optimization PC-DARTS [49] 74.9 92.2 0.1 5.3 Gradient-based PNASNet-5[23] 74.2 91.9 45 5.1 Model-based Predictor GHN[53] 73.0 91.3 0.84 5.7 Model-based Predictor NAONet[27] 74.3 91.8 200 11.35 Model-based Predictor TG-NAS (ours) 74.9 92.2 0.0014 5.6 Model-based Predictor *All results are searched on CIFAR-10 and evaluated on ImageNet if not specialized noted
Notably, despite the efficiency claims of zero-cost NAS methods over conventional NAS due to avoiding the training of sampled architectures, variations in computational cost and search time persist among them. For example, TE-NAS [3] required over 4 GPU hours, while ZiCo [17] demanded over 10 GPU hours for the search on ImageNet on the DARTS space. Despite the GPU differences, this discrepancy arises from the fact that several proxy computations necessitate at least one forward pass or a forward-backward pass, requiring multiple runs for result stabilization. Additionally, many are data-dependent, limiting their general applicability and extending search times.
In contrast, TG-NAS completed the search in about 40 seconds, achieving up to a 300×300$\times$300 × speedup compared to other zero-cost methods due to the lightweight nature of the proposed predictor model. Notably, the comparison of search times does not factor in the predictor model training time (the model training required 1.5 GPU hours), as it is a one-time training effort, and the pre-trained model can be distributed for use in various scenarios and search spaces.
4.4 NAS Result for ImageNet on the DARTS Search Space
For the DARTS search space, the final discovered cell architecture is shown in the Appendix. After the target cell was determined, the final network was constructed by stacking 14 cells, with the initial channel number set to 48. Performance results of TG-NAS discovered architectures compared against previous NAS works are shown in Table 7. Our approach achieves a top-1/5 accuracy of 74.9% and 92.2% respectively. Notably, TG-NAS achieves up to 300$\times$ improvement in search efficiency compared to previous state-of-the-art zero-cost proxy methods. Furthermore, in comparison to other predictor-based NAS approaches, TG-NAS achieves over a 600$\times$ increase in search efficiency, completing the search process in less than two minutes on a single NVIDIA RTX 4090 GPU.
@@ 5 Conclusion
In this work, we propose TG-NAS, a model-based zero-cost (ZC) proxy that is broadly applicable to new search spaces containing previously unseen operators. TG-NAS integrates text-based operator descriptions—processed by a fine-tuned sentence transformer—with a graph convolutional network (GCN) predictor, enabling it to function effectively as a ZC proxy. It offers key advantages in robustness, generalizability, proxy independence, and cost-effectiveness. Our experiments demonstrate that TG-NAS consistently outperforms existing proxies across a wide range of NAS benchmarks, establishing it as a strong foundational component for efficient architecture search. TG-NAS achieves up to 300× improvement in search efficiency over prior state-of-the-art ZC methods. Notably, it discover competitive models with 93.75% CIFAR-10 accuracy on the NAS-Bench-201 space and 74.9% ImageNet top-1 accuracy on the DARTS space.